If Sin x +sin y=1/3 Cos x+cos y=1/4 What is sin(x+y)

If 
Sin x +sin y=1/3
Cos x+cos y=1/4
What is sin(x+y)

Grade:11

1 Answers

Arun
5 years ago
Dear student
 
2 sin x+y/2 cos x-y/2 = 1/3
&
2 cos x+y/2 cos x-y/2 = ¼
 
Now divide
tan (x+y/2) = 4/3
 
Now
 
Sin(x+y) = 2 tan (x+y/2) /1 + tan² (x+y/2)
 
 = (8/3) /(1 + 16/9)
 
 = 24/25